I am suppose to find the slope-intercept form ofthe equation of the line that passes through the points.
Points (-8,1) (-8,7)
This is what I did.
m=(7-1)/(-8--8)
=-6/0
So the slop is undefined or vertical. So I assumed the answer was x=6. However, the correct answer was x+8=0. How is that slope intercept form? And how did you arrive at that solution. Please help. It would be greatly appreciated!

Answer by Alan3354(69443)

I am suppose to find the slope-intercept form ofthe equation of the line that passes through the points.
Points (-8,1) (-8,7)
This is what I did.
m=(7-1)/(-8--8)
=-6/0
So the slop is undefined or vertical. So I assumed the answer was x=6. However, the correct answer was x+8=0. How is that slope intercept form? And how did you arrive at that solution. Please help. It would be greatly appreciated!
---------
Since the slope is undefined, there's no y in the equation. There's no way to make it y = mx + b.
---------
You can see from these 2 points that for all values of y, x = -8.
The equation can only be x = -8, or x+8 = 0.